You found the following stock quote for DRK Enterprises, Inc., at your favorite website. You also found that the stock paid an annual dividend of $1.32, which resulted in a dividend yield of 1.7 percent. Assume the company has 71 million shares of stock outstanding and a P/E ratio of 16.

DAILY YTD
  Company Symbol Vol Close Chg %Chg %Chg
  DRK Enterprises DRK 18,649,130 ?? .26 .45% 8.73%

What was net income for the most recent four quarters? (Do not round intermediate calculations. Round your answer to the nearest whole number.)
Answer #1

Dividend yield = .017

P0 = $1.32/ .017= $77.64

Stock closed up $0.26, so yesterday’s closing price = $77.64 – 0.26 = $77.38

PE = 16

EPS = P0 /16

= 77.64/16

= $4.8525

EPS = Net Income/ Shares

Net Income = Shares x EPS

= 71,000,000 * 4.8525

= $ 344,527,500